OpenStudy (dottiecat):

A TV that usually sells for $182.19 is on sale for 15% off. If sales tax on the TV is 5%, what is the price of the TV, including tax? A $164.15 B $217.26 C $161.06 D $162.60

OpenStudy (dottiecat):

@escamer @DIDITHEDAD

OpenStudy (dottiecat):

@joojoe

OpenStudy (escamer):

D. 182.19 * .15 = 27.32 182.19 - 27.32 = 154.86 * .05 = 7.713 + 154.86 = 162.60 :)

OpenStudy (dottiecat):

Thanks!! Can you help me with 1 more?

OpenStudy (escamer):

Sure, just tag me in the question.

OpenStudy (dottiecat):

Thanks

OpenStudy (escamer):

Anytime